OpenStudy (princess):

how to get the gcf of 16 & 39

OpenStudy (anonymous):

16 = 1*2^4 39 = 1*3*13 gcd are the products that are common, so gcd (16,39) = 1. In fact if two numbers have the property gcd(a,b) =1, a and b are said to be "relatively prime"

OpenStudy (anonymous):

factors of 16 are 1,2, 4, 8,16 right. also look factors of 39 are 1,3,13,and 39 the common factor is 1 which is the highest ok. then gcd=1

OpenStudy (anonymous):

More commonly we write out a number as a product of prime factors as this makes intersection of prime factors obvious.